A person crosses a 600m long street in 5 minutes, what is his speed in km per hour?
A. 3.6
B. 7.2
C. 8.4
D. 10

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this Class VI Maths answer. We will update the answer very soon.
Janarth
one year ago
If the length of street = 600m,
i am going to convert to KM which is 0.6 km

time is 5 mins which is 1 / 12 hours

Using the formula  Speed = \frac{distance}{time},

\frac{0.6}{\frac{1}{12}} = 7.2 Km/Hr

so the answer is b